The time in hours a transistor lasts follows an exponential distribution with a mean of 100. Find the probability that a transistor lasts longer than

a) 15 hours.

b) 110 hours

c) 110 hours given that it lasts longer than 95 hours.



1
Expert's answer
2021-06-20T18:42:29-0400

Let "X=" the time a transistor lasts : "X\sim Exp(\lambda)"

Given "\lambda=\dfrac{1}{\mu}=\dfrac{1}{100}=0.01"

a)


"P(X>15)=e^{-0.01(15)}\approx0.860708"

b)


"P(X>110)=e^{-0.01(110)}\approx0.332871"

c)


"P(X>110|X>95)=\dfrac{P((X>110)\cap (X>95))}{P(X>95)}"

"=\dfrac{P(X>110)}{P(X>95)}=\dfrac{e^{-0.01(110)}}{e^{-0.01(95)}}=e^{-0.01(15)}"

"=P(X>15)\approx0.860708"